# GP-GPT: Large Language Model for Gene-Phenotype Mapping

### v3

@@ 2.5 Gene Networks and Knowledge Graphs
Whole genome sequencing [67, 68] and GWAS studies [69, 70]have already provided substantial evidence for investigating disease-associated genes. Recent research suggests that genomics datasets can be represented as gene networks, where the nodes symbolize genes, phenotypes, or proteins, and the connections between the nodes indicate the relationships between these entities. Schlitt, et.al.  [71] propose a novel method for identifying functionally related genes by comparing neighbourhoods within gene networks. This approach does not depend on gene sequence or protein structure homologies and can be applied to any organism and a wide range of experimental datasets. Gene networks enable researchers to seek additional evidence to cross-check their results using various sources, e.g. enriching existing computational approaches such as linkage analysis and GWAS studies by including the relationships identifications between genes from protein-protein interaction (PPI), gene expression, gene ontology and other domain knowledge  [72]. HumanNet, HumanNet2, and HumanNet3 constructed genome-scale functional networks for human genes using a variety of evidence sources, including human mRNA co-expression, protein-protein interactions, protein complexes, and comparative genomics datasets. In these networks, the edges may represent different types of associations between genes, which can be mapped through both experimental and computational methods [73, 7, 74].
At the same time, the development of knowledge bases and graphs has greatly enhanced performance in fields like computer vision and natural language processing. Likewise, knowledge graphs have been widely applied to specific biological tasks, such as cancer diagnosis [75, 76, 77]. Other studies take advantage of the Gene Ontology (GO) knowledgebase to build the neural network architecture to simulate the gene or protein reactions in the cells [76, 77]. Combining multiple bio-factors, biological factor regulatory neural network (BFReg-NN)  [78] aims at simulating the complex biological processes in a cell system, understanding the functions of genes or proteins.
@@ 3.1 Data
The datasets in genomics, proteomics, and medical genetics have experienced significant growth in both scale and diversity. Our dataset-building attempt starts from the gene entities and their textual descriptions maintained by the National Center for Biotechnology Information (NCBI) [79]. Taking the 192064 individual gene items from NCBI as the entries, we then collected and screened the data of related bio-factors from multiple datasets. Each of these datasets stands for a particular level in the genomics system(Fig. 1).
@@
For the purpose of comprehensively incorporating and extracting the mapping between genes and associated phenotype/diseases, multiple data sources were integrated into the training dataset of this study. A major part of the dataset came from the Online Mendelian Inheritance in Man(OMIM) database [80], which is a comprehensive and authoritative resource for understanding human genes and their associated genetic disorders. OMIM keeps full-text information on all known Mendelian disorders and over 16,000 genes and provides detailed information on human genes, genetic phenotypes, and the relationships between genes and phenotypes. Genes often cast influences on phenotypes through gene products and a series of related complex biological pathway alterations  [8]. Building on this concept, we incorporated the information of gene products from UniPort [81], and their molecular function information from NCBI. The extracted pairs of genes and associated phenotypes/diseases were later verified and supplied through dbGaP [26] and DisGeNet [9]. In summary, from OMIM, we collected the gene-phenotype pairs along with the relational graphs. A total of 4401 pairs of phenotype and gene have been collected and verified with the dbGaP data. From UniPort, we collected 1286 gene-protein pairs. From DisGeNet, we use the Relation Extraction (RE) samples from the evidence terms. We integrated these multi-level genomics text data to serve our training objective. All the mentioned data are examined and matched to the NCBI gene IDs and organized accordingly.
@@ 3.2 Model
In this study, we utilize the parameter-efficient fine-tuning techniques of the LLaMA model family (Llama2, Llama3, Llama3.1) using Low-Rank Adaptation (LoRA) [39], and QLoRA [42] technique. The objective is to adapt the Llama3 model for specific downstream tasks while minimizing the number of trainable parameters and computational resources required. LoRA introduces a low-rank adaptation matrix into the transformer architecture, allowing efficient fine-tuning with fewer parameters. In addition, QLoRA applies a quantized pre-trained language model into low-rank adapters, significantly reducing memory usage to further finetune large language models. Both techniques modify the original model minimally, making it computationally efficient and less prone to overfitting, especially with limited data. In our study, we tested both LoRA and 8-bit QLoRA on the Llama2 7B model and the Llama3.1 8B model. With no significant performance discrepancies between the two techniques, we employed 8-bit QLoRA to achieve efficient fine-tuning. Additionally, 4-bit QloRA was used during the fine-tuning of the Llama3.1 70B model.
@@
Fine-tuning: The entire instruction fine-tuning process can be considered as a two-stage training approach. In the first stage, the training content is drawn from the initial segment of the training set, which is prepared for fine-tuning through instruction mask prediction. In the second stage, the training content is taken from the latter part of the training set, which is structured for supervised fine-tuning using a question-answer format. GP-GPT can be fine-tuned on various downstream tasks involving molecules, proteins, and text. In both stages, to bridge the gap between pre-training and fine-tuning stages and unify different genomic downstream tasks, we employ a prompt-based fine-tuning approach that converts various task formats into a sequence generation format [82, 83].
@@ 3.3 Evaluation
Case studies: We first evaluate GP-GPT on the text generation task. The models were required to answer the questions about gene and disease knowledge. The ’Gene disease association’ questions from the QA database GeneTuring [84] have been used to test our model, and an expert evaluation has been conducted by comparing the model-generated texts and golden standard answers.
@@
We evaluate the GP-GPT against other 7 large language models. The setting of the same prompt components has been used across all different models. The candidate LLM includes: original Llama2(7B), Llama2(13B), Llama3(8B), Llama3.1(8B), Llama3.1(70B), and the Bio-GPT [53], and the GPT-4 model [11]. Our evaluation of GP-GPT is more rigorous than the evaluation of LLMs and the GPT models in genomics [84], which minimized manual involvement and emphasized both exact matches and partial correctness.
@@
To evaluate the model’s understanding of the relation between genes and phenotype/diseases, the well-established relation determination task has been applied to directly test the model’s performance. The relation determination task was formulated as a text generation task with the proper question-answer prompt. The relation determination prompt is composed of instruction and text, models are asked to generate answers in yes/no based on the text. We use the test set from TBGA [32] data. We screened the testing data by checking the gene-disease pairs with their corresponding pairs in DisGeNet data, to ensure the knowledge correctness and testing effectiveness. We tested the GP-GPT with other large language models, for the local running models, the results were collected in a ’mixed-prompt’ strategy. Thus, for GP-GPT, Llama2, Llama3, Llama3.1, and Bio-GPT, each single question-answer testing data is formulated with multiple prompts. During the results collection, the best answer was selected as the final output of the model for further metric calculations. But for commercial large language models like GPT-4, only one kind of prompt was tested due to its cost.
@@ 5.1 Limitations
The data used in this study covers multiple data sources, including dbGaP, UniPort, DisGeNet, NCBI, and OMIM. Among these datasets, the employed bio-text data are mainly based on OMIM. Although the training corpus is composed of 4401 pairs of phenotype and gene, the construction of training contextual data is heavily dependent on high-quality textual data, which is lacking for most genes and diseases. With this limitation of data sources, there is demand of generating more reliable genomics textual data from other structure dataset. Inspired by the network-based disease gene prediction research, using the network data of protein-protein interaction (PPI), gene expression, gene ontology, and other domain knowledge to generate large relational-genomics text data is an available way. Based on conventional relationships often observed in bio-networks, e.g. dominating, inhibiting, and promoting, the relation network of bio-factors can provide rich text data following a well-designed graph-to-text interpreting system. To be noted that, such systems are usually based on LLMs as well [85, 86, 87].
@@
Tokenization is one of the essential steps in the success of the training language model. To simplify the model training, we adopt the character-based tokenization method from the original Llama2, and Llama3 for all the genomics entities. One reason for this is that the model is only focusing on the name-entities features and relation description, which both are based on natural languages. Another reason is that Llama2 and Llama3 employ a character-based method for the numerical tokenization approach, which has been shown to be more effective across various arithmetic tasks. In practice, these tokenizers work well for the current stage of GP-GPT. However, a from-scratch pre-training can be necessary for a better representation of genetics, genomics, and other levels of bio-factors, especially when more features from the sequence are involved. Previous works [43, 44, 46] have provided available instances for bioinformatics costumed tokenizor. Due to the computation cost, we decide to put such endeavours into our further improvement of GP-GPT.
@@
Furthermore, the evaluation methods for such genomics language models are still underdeveloped. The direct issue to this end is that the missing of standard benchmarks in this area. On one hand, traditional metrics such as BLEU [88], and ROUGE [89] do not fit the defined research aim in this study, thus the metrics lack the necessary focus on certain bio-factor entities. In this study, we use modified BLEU which adjusts the weighting for genomics keywords to mediate this issue. Both the original BLEU score and modified BLEU score are reported for GP-GPT, leading to a more comprehensive estate of model performance. On the other hand, a stand-alone test set for evaluation becomes a problem for this study. That is due to the aim of the model being to build a fundamental language AI system to interpret real-world genetic medical relations. Our solution in the study is based on internal validation on a random select genomics relationship which has been verified by biological experiments.
@@ 5.2 Future Perspectives
Our proposed large language models are not grounded in incorporating other domains of data, such as biological sequence data and medical images. The wide applicability of transformer architecture [90] facilitates modality fusion. Recent research in computer vision and NLP also demonstrates significant benefits in combining multiple modalities to achieve application success [91, 92, 93, 94, 95, 96]. Domain-specific models are natural platforms for incorporating data from other modality to improve and scale practical usability [97, 98, 99, 100, 101, 102, 103, 104, 105, 106].
Bio-sequence data counts for the most dominant part of all available data in bioinformatics studies. One of the efficient ways to combine the bio-sequence data and LLMs is utilizing the parallel architecture to place both the sequence-specific model and language-specific model [107]. Using specific tokenizers to encode particular sequence features is another practical method [108]. In future versions of GP-GPT, the sequence data from DNA sequences, gene variances, and gene products(small-RNA, protein) will be included in the framework.
Imaging genetics offers the possibility of building connections between genetics insights into the biology of complex diseases and functions [109, 110, 111]. LLM systems can largely benefit the representation of gene variants and candidate phenotypes, thus providing a qualitative description of domain knowledge. By utilizing the image encoder, bio-factor entity encoder, and variants encoder, this framework can capture the intricate interplay in semantic embedding space, which leads to more comprehensive association identifications.

### v4

@@ 2.5 Gene Networks and Knowledge Graphs
Whole genome sequencing [cho2016muffinn, horn2018netsig] and GWAS studies [greene2015understanding, shim2017gwab]have already provided substantial evidence for investigating disease-associated genes. Recent research suggests that genomics datasets can be represented as gene networks, where the nodes symbolize genes, phenotypes, or proteins, and the connections between the nodes indicate the relationships between these entities. Schlitt, et.al. [schlitt2003gene] propose a novel method for identifying functionally related genes by comparing neighbourhoods within gene networks. This approach does not depend on gene sequence or protein structure homologies and can be applied to any organism and a wide range of experimental datasets. Gene networks enable researchers to seek additional evidence to cross-check their results using various sources, e.g. enriching existing computational approaches such as linkage analysis and GWAS studies by including the relationships identifications between genes from protein-protein interaction (PPI), gene expression, gene ontology and other domain knowledge [ata2021recent]. HumanNet, HumanNet2, and HumanNet3 constructed genome-scale functional networks for human genes using a variety of evidence sources, including human mRNA co-expression, protein-protein interactions, protein complexes, and comparative genomics datasets. In these networks, the edges may represent different types of associations between genes, which can be mapped through both experimental and computational methods [kulikajevas2021humannet, hwang2019humannet, kim2022humannet].
At the same time, the development of knowledge bases and graphs has greatly enhanced performance in fields like computer vision and natural language processing. Likewise, knowledge graphs have been widely applied to specific biological tasks, such as cancer diagnosis [elmarakeby2021biologically, yu2016translation, ma2018using]. Other studies take advantage of the Gene Ontology (GO) knowledgebase to build the neural network architecture to simulate the gene or protein reactions in the cells [yu2016translation, ma2018using]. Combining multiple bio-factors, biological factor regulatory neural network (BFReg-NN) [dai2023biological] aims at simulating the complex biological processes in a cell system, understanding the functions of genes or proteins.
@@ 3.1 Data
The datasets in genomics, proteomics, and medical genetics have experienced significant growth in both scale and diversity. Our dataset-building attempt starts from the gene entities and their textual descriptions maintained by the National Center for Biotechnology Information (NCBI) [ncbi_url]. Taking the 192064 individual gene items from NCBI as the entries, we then collected and screened the data of related bio-factors from multiple datasets. Each of these datasets stands for a particular level in the genomics system(Fig. 1).
@@
For the purpose of comprehensively incorporating and extracting the mapping between genes and associated phenotype/diseases, multiple data sources were integrated into the training dataset of this study. A major part of the dataset came from the Online Mendelian Inheritance in Man(OMIM) database [omim_url], which is a comprehensive and authoritative resource for understanding human genes and their associated genetic disorders. OMIM keeps full-text information on all known Mendelian disorders and over 16,000 genes and provides detailed information on human genes, genetic phenotypes, and the relationships between genes and phenotypes. Genes often cast influences on phenotypes through gene products and a series of related complex biological pathway alterations [amberger2015omim]. Building on this concept, we incorporated the information of gene products from UniPort [coudert2023annotation], and their molecular function information from NCBI. The extracted pairs of genes and associated phenotypes/diseases were later verified and supplied through dbGaP [mailman2007ncbi] and DisGeNet [pinero2020disgenet]. In summary, from OMIM, we collected the gene-phenotype pairs along with the relational graphs. A total of 4401 pairs of phenotype and gene have been collected and verified with the dbGaP data. From UniPort, we collected 1286 gene-protein pairs. From DisGeNet, we use the Relation Extraction (RE) samples from the evidence terms. We integrated these multi-level genomics text data to serve our training objective. All the mentioned data are examined and matched to the NCBI gene IDs and organized accordingly.
@@ 3.2 Model
In this study, we utilize the parameter-efficient fine-tuning techniques of the LLaMA model family (Llama2, Llama3, Llama3.1) using Low-Rank Adaptation (LoRA) [hu2021lora], and QLoRA [dettmers2024qlora] technique. The objective is to adapt the Llama3 model for specific downstream tasks while minimizing the number of trainable parameters and computational resources required. LoRA introduces a low-rank adaptation matrix into the transformer architecture, allowing efficient fine-tuning with fewer parameters. In addition, QLoRA applies a quantized pre-trained language model into low-rank adapters, significantly reducing memory usage to further finetune large language models. Both techniques modify the original model minimally, making it computationally efficient and less prone to overfitting, especially with limited data. In our study, we tested both LoRA and 8-bit QLoRA on the Llama2 7B model and the Llama3.1 8B model. With no significant performance discrepancies between the two techniques, we employed 8-bit QLoRA to achieve efficient fine-tuning. Additionally, 4-bit QloRA was used during the fine-tuning of the Llama3.1 70B model.
@@
Fine-tuning: The entire instruction fine-tuning process can be considered as a two-stage training approach. In the first stage, the training content is drawn from the initial segment of the training set, which is prepared for fine-tuning through instruction mask prediction. In the second stage, the training content is taken from the latter part of the training set, which is structured for supervised fine-tuning using a question-answer format. GP-GPT can be fine-tuned on various downstream tasks involving molecules, proteins, and text. In both stages, to bridge the gap between pre-training and fine-tuning stages and unify different genomic downstream tasks, we employ a prompt-based fine-tuning approach that converts various task formats into a sequence generation format [brown2020language, gaoetal2021making].
@@ 3.3 Evaluation
Case studies: We first evaluate GP-GPT on the text generation task. The models were required to answer the questions about gene and disease knowledge. The ’Gene disease association’ questions from the QA database GeneTuring [hou2023geneturing] have been used to test our model, and an expert evaluation has been conducted by comparing the model-generated texts and golden standard answers.
@@
We evaluate the GP-GPT against other 7 large language models. The setting of the same prompt components has been used across all different models. The candidate LLM includes: original Llama2(7B), Llama2(13B), Llama3(8B), Llama3.1(8B), Llama3.1(70B), and the Bio-GPT [luo2022biogpt], and the GPT-4 model [achiam2023gpt]. Our evaluation of GP-GPT is more rigorous than the evaluation of LLMs and the GPT models in genomics [hou2023geneturing], which minimized manual involvement and emphasized both exact matches and partial correctness.
@@
To evaluate the model’s understanding of the relation between genes and phenotype/diseases, the well-established relation determination task has been applied to directly test the model’s performance. The relation determination task was formulated as a text generation task with the proper question-answer prompt. The relation determination prompt is composed of instruction and text, models are asked to generate answers in yes/no based on the text. We use the test set from TBGA [marchesin2022tbga] data. We screened the testing data by checking the gene-disease pairs with their corresponding pairs in DisGeNet data, to ensure the knowledge correctness and testing effectiveness. We tested the GP-GPT with other large language models, for the local running models, the results were collected in a ’mixed-prompt’ strategy. Thus, for GP-GPT, Llama2, Llama3, Llama3.1, and Bio-GPT, each single question-answer testing data is formulated with multiple prompts. During the results collection, the best answer was selected as the final output of the model for further metric calculations. But for commercial large language models like GPT-4, only one kind of prompt was tested due to its cost.
@@ 5.1 Limitations
The data used in this study covers multiple data sources, including dbGaP, UniPort, DisGeNet, NCBI, and OMIM. Among these datasets, the employed bio-text data are mainly based on OMIM. Although the training corpus is composed of 4401 pairs of phenotype and gene, the construction of training contextual data is heavily dependent on high-quality textual data, which is lacking for most genes and diseases. With this limitation of data sources, there is demand of generating more reliable genomics textual data from other structure dataset. Inspired by the network-based disease gene prediction research, using the network data of protein-protein interaction (PPI), gene expression, gene ontology, and other domain knowledge to generate large relational-genomics text data is an available way. Based on conventional relationships often observed in bio-networks, e.g. dominating, inhibiting, and promoting, the relation network of bio-factors can provide rich text data following a well-designed graph-to-text interpreting system. To be noted that, such systems are usually based on LLMs as well [chai2023graphllm, swamy2021interpreting, hu2023beyond].
@@
Tokenization is one of the essential steps in the success of the training language model. To simplify the model training, we adopt the character-based tokenization method from the original Llama2, and Llama3 for all the genomics entities. One reason for this is that the model is only focusing on the name-entities features and relation description, which both are based on natural languages. Another reason is that Llama2 and Llama3 employ a character-based method for the numerical tokenization approach, which has been shown to be more effective across various arithmetic tasks. In practice, these tokenizers work well for the current stage of GP-GPT. However, a from-scratch pre-training can be necessary for a better representation of genetics, genomics, and other levels of bio-factors, especially when more features from the sequence are involved. Previous works [bryant2022improved, abramson2024accurate, pei2023biot5] have provided available instances for bioinformatics costumed tokenizor. Due to the computation cost, we decide to put such endeavours into our further improvement of GP-GPT.
@@
Furthermore, the evaluation methods for such genomics language models are still underdeveloped. The direct issue to this end is that the missing of standard benchmarks in this area. On one hand, traditional metrics such as BLEU [papineni2002bleu], and ROUGE [lin2004rouge] do not fit the defined research aim in this study, thus the metrics lack the necessary focus on certain bio-factor entities. In this study, we use modified BLEU which adjusts the weighting for genomics keywords to mediate this issue. Both the original BLEU score and modified BLEU score are reported for GP-GPT, leading to a more comprehensive estate of model performance. On the other hand, a stand-alone test set for evaluation becomes a problem for this study. That is due to the aim of the model being to build a fundamental language AI system to interpret real-world genetic medical relations. Our solution in the study is based on internal validation on a random select genomics relationship which has been verified by biological experiments.
@@ 5.2 Future Perspectives
Our proposed large language models are not grounded in incorporating other domains of data, such as biological sequence data and medical images. The wide applicability of transformer architecture [vaswani2017attention] facilitates modality fusion. Recent research in computer vision and NLP also demonstrates significant benefits in combining multiple modalities to achieve application success [radford2021learning, liu2024sora, team2024chameleon, lee2023multimodality, dou2023towards, xiao2023instruction]. Domain-specific models are natural platforms for incorporating data from other modality to improve and scale practical usability [liu2024understanding, liu2024surviving, li2023artificial, wang2024large, zhao2022embedding, liu2023radonc, cai2023exploring, cai2023multimodal, liu2023tailoring, tan2023promises].
Bio-sequence data counts for the most dominant part of all available data in bioinformatics studies. One of the efficient ways to combine the bio-sequence data and LLMs is utilizing the parallel architecture to place both the sequence-specific model and language-specific model [xu2023protst]. Using specific tokenizers to encode particular sequence features is another practical method [heinzinger2023bilingual]. In future versions of GP-GPT, the sequence data from DNA sequences, gene variances, and gene products(small-RNA, protein) will be included in the framework.
Imaging genetics offers the possibility of building connections between genetics insights into the biology of complex diseases and functions [bigos2010imaging, bi2023community, bi2023structure]. LLM systems can largely benefit the representation of gene variants and candidate phenotypes, thus providing a qualitative description of domain knowledge. By utilizing the image encoder, bio-factor entity encoder, and variants encoder, this framework can capture the intricate interplay in semantic embedding space, which leads to more comprehensive association identifications.